A project has annual cash flows of $4,500 for the next 10 years and then $9,500 each year for the following 10 years. The IRR of this 20-year project is 13.04%. If the firm's WACC is 10%, what is the project's NPV? Round your answer to the nearest cent.
Answer #1

Project's NPV is $ 10,668.89

Step-1:Calculation of cost of project
IRR is the rate at which present value of cash inflows is equal to its cost.
Present value of annuity of 1 for 10 years = (1-(1+i)^-n)/i Where,
= (1-(1+0.1304)^-10)/0.1304 i = 13.04%
= 5.41757987 n = 10
Present value of 1 = (1+i)^-n
= (1+0.1304)^-10
= 0.29354758
Present value of cash flows of first 10 years = $   4,500.00 * 5.41758 = $ 24,379.11
Present value of cash flows of next 10 years = $   9,500.00 * 5.41758 * 0.293548 = $ 15,108.02
Present value of cash flows of next 20 years $ 39,487.13
So, cost of project is $ 39,487.13
Step-2:Calculation of present value of project's cash inflows at WACC that is discounted rate
Present value of annuity of 1 for 10 years = (1-(1+i)^-n)/i Where,
= (1-(1+0.10)^-10)/0.10 i = 10.00%
= 6.14456711 n = 10
Present value of 1 = (1+i)^-n
= (1+0.10)^-10
= 0.38554329
Present value of cash flows of first 10 years = $   4,500.00 * 6.144567 = $ 27,650.55
Present value of cash flows of next 10 years = $   9,500.00 * 6.144567 * 0.385543 = $ 22,505.47
Present value of cash flows of next 20 years $ 50,156.02
Step-3:Calculation of project's NPV
Present value of cash inflow $ 50,156.02
Less cost of project $ 39,487.13
Project's NPV $ 10,668.89